OpenStudy (anonymous):

write an eqaution of the line containg the given point and perpendicular to the given line. (0,8); 2x+7y=5 the equation of the line is y= ?

OpenStudy (abmon98):

7y=-2x+5 y=-2/7x+5 m=-1/-2/7 m=7/2 y-8=7/2(x-0) 2y-16=7x y=7/2x+16